# Reasoning with Confidence: Efficient Verification of LLM Reasoning Steps via Uncertainty Heads

### v1

@@ Appendix A Additional Experimental Results
To verify our framework works for models of different size, family, and post-training, we conduct step-level correctness and offline BoN evaluation on a Phi-4 UHead trained on Qwen3-8B annotated training data. The step-level correctness prediction and Offline BoN results are presented in Appendix A and Appendix A correspondingly. For budget reason, we do not perform DeepSeek-R1 annotation for Phi-4 training data. On step-level correctness, Qwen3-8B-annotated UHead achieves the best performance in Meeting and Calendar Planning, and the best average performance on OOD tasks. On StrategyQA and overall average, it ranks the second.
For offline BoN, the Qwen3-8B–annotated UHead outperforms two strong Qwen2.5-Math PRMs on MATH and matches the strongest PRMs on GSM8K. On ProofNet and StrategyQA, it also outperforms some much larger PRMs.
Table 6: Beam-Search decoding accuracy across datasets (Qwen3-8B).
Method # Sample Math (ID) Planning (OOD) QA (OOD) Average
[…]
Qwen2.5-Math-7B-PRM800k 263K 88.6
Qwen2.5-Math-PRM-7B 860K 78.4 87.8
[…]
UHead Self-anno (Ours) 32K 76.0 87.8
UHead DeepSeek-anno (Ours) 32K 75.6 89.0
Table 7: PR-AUC for detecting incorrect reasoning steps for Phi-4. Best scores are shown in bold, and other competitive scores are underlined. # Sample indicates the number of training samples, where each sample corresponds to a reasoning trajectory with step-level annotations.  Qwen2.5-Math-PRM-7B’s training data is filtered from an 860K-sample dataset; the exact size after filtering is not specified in their paper.
@@
[…]
UHead Qwen3-8B-anno (Ours) 32K .404 .340 .155 .756 .646 .592 .347 .347 .300 .538 .448
Table 8: Offline Best-of-NN decoding accuracy across datasets for Phi-4 model. For datasets with verifiable final answer, we also provide Majority Voting. We use max uncertainty aggregation over steps.
@@
UHead Qwen3-8B-anno (Ours) 32K 95.9 100. 93.1 93.0 95.7
Table 9: PR-AUC performance of UHeads trained on the most Diverse/Indiverse 2K questions (3 trajectories sampled for each question). Data diversity benefits the overall performance of UHead.
@@
UHead DeepSeek-anno Indiverse 6K .308 .549 .205 .626 .687 .685 .377 .251 .354 .525 .461
UHead DeepSeek-anno Diverse 6K .409 .575 .180 .707 .793 .792 .271 .325 .388 .578 .507
Figure 3: Best-of-NN performance on GSM8k and ScienceQA for different $N$ values (Qwen3-8B).
@@
Figure 4: Top row: PR-AUC of UHeads with increasing training set size (x-axis). Bottom row: scaling training data either by adding new unique questions or by sampling additional trajectories. One dataset per domain is shown.
Appendix B Discussion and Limitations
§ 5 shows that UHead performance increases with larger training data and benefits from data diversity. Curves in Figure 2 and Figure 4, although show diminishing marginal gains, do not seem to reach the top for tasks like StrategyQA. Therefore, it seems possible to further unleash the potential of UHead with further data scaling – sampling more reasoning trajectories per question or adding new questions beyond the PRM800K training set. In this work, we do not involve questions outside the PRM800K training set to establish a fair comparison with PRMs trained on data derived from this set (e.g., Qwen2.5-Math-7B-PRM800K). Due to the high cost of DeepSeek-R1 (annotation of 32K reasoning trajectories cost >2000{>}2000 USD), we also do not annotate more reasoning trajectories per question. We leave this promising scaling to future work, including the integration of diverse, high-quality questions outside the math domain. To save annotation cost, future work may also substitute API-based LLM with recent small but capable reasoning models, for example, GPT-OSS (agarwal2025gptoss).
Another limitation regarding the UHead framework is that it needs to be trained on the internal signal of target LLMs it supervises. Once trained, a UHead cannot be directly applied to another LLM, since it depends on model-specific internal states. By contrast, PRMs can supervise any LLM because they only operate on output tokens. However, UHeads are highly efficient in parameter size and training data, making their training relatively inexpensive. Once trained, UHeads can significantly save inference costs compared to PRMs. In practice, applications may focus on a small number of widely used models. If UHeads for these models are shared online (e.g., on HuggingFace), practitioners could readily download and apply them without training from scratch. Future work may also investigate fine-tuning UHeads to adapt to customized or fine-tuned versions of target LLMs.
Appendix C Prompts
C.1 LLM Prompt

### v3

@@ Appendix A Additional Experimental Results
A.1 Results with Phi-4
To verify that our framework works for models of different sizes, families, and post-training approaches, we conduct step-level correctness and BoN evaluation on a Phi-4 ReProbe trained on Qwen3-8B annotated training data. The step-level correctness prediction and BoN results are presented in Table 6 and Table 7 correspondingly. On step-level correctness, ReProbe achieves the best performance in Meeting and Calendar Planning, and the best average performance on OOD tasks. On StrategyQA and overall average, it ranks second.
For BoN, the Qwen3-8B–annotated ReProbe outperforms two strong Qwen2.5-Math PRMs on MATH and matches the strongest PRMs on GSM8K. On ProofNet and StrategyQA, it also outperforms some much larger PRMs.
A.2 ReProbe under Native Thinking Mode
Native LRM reasoning often occurs within <think> tags without explicit step separations. To evaluate ReProbe in this setting, we train a ReProbe model for Qwen3-1.7B in native thinking mode. We modify the experimental setup from § 4 in two ways: (1) Since free-form reasoning from LRMs is much longer than formatted CoT, we replace the DeepSeek-R1 API with a locally deployed GPT-OSS-120B for step-level correctness annotation to reduce API costs. This model outperforms the January version of DeepSeek-R1 on multiple benchmarks (e.g., GPQA Diamond, AIME 2024). (2) Without format-based step separations, we treat each sentence as a step and prompt GPT-OSS-120B to identify errors conditioned on all previous sentences (prompts in Appendix F). Results in Table 8 show that ReProbe outperforms PRMs on ID and OOD average scores in native thinking mode.
A.3 Ablation Study on Data Quantity and ReProbe Architectures
Table 9 shows that the linear probe and token-level ReProbe are worse than the proposed step-level ReProbe. Fig 5 shows the per-dataset performance improvement by scaling the data quantity.
@@ A.3 Ablation Study on Data Quantity and ReProbe Architectures
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Attn+Logit, Qwen3-8B-anno (Ours) 32K .404 .340 .155 .756 .646 .592 .347 .347 .300 .538 .448
Table 6: PR-AUC for detecting incorrect reasoning steps for Phi-4. Best scores are shown in bold, and other competitive scores are underlined. # Sample indicates the number of training samples, where each sample corresponds to a reasoning trajectory with step-level annotations. $\ddagger$ Qwen2.5-Math-PRM-7B’s training data is filtered from an 860K-sample dataset; the exact size after filtering is not specified in their paper.
@@
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Attn+Logit, Qwen3-8B-anno (Ours) 32K 95.9 100. 93.1 93.0 95.7
Table 7: Best-of-NN decoding accuracy across datasets for Phi-4 model. For datasets with verifiable final answer, we also provide Majority Voting.
Method # Sample Math (ID) Planning (OOD) QA (OOD) Average
@@
Math-Shepherd-PRM-7B 440K .049 .248 .234 .219 .481 .338 .265 .313 .177 .323 .268
RLHFlow-PRM-Deepseek-Data 253K .063 .267 .228 .169 .498 .322 .449 .368 .186 .361 .296
RLHFlow-PRM-Mistral-Data 273K .047 .184 .164 .141 .457 .275 .273 .279 .132 .285 .228
Universal-PRM-Qwen2.5-Math-7B 690K .175 .331 .299 .272 .708 .446 .335 .329 .268 .418 .362
Qwen2.5-Math-7B-PRM800K 263K .217 .392 .505 .270 .716 .587 .391 .336 .371 .460 .427
Qwen2.5-Math-PRM-7B 860K .293 .270 .465 .328 .732 .605 .378 .293 .343 .467 .421
Reasoning Probes (ReProbes)
ReProbe, Hidden States, GPT-OSS-anno (Ours) 10.8K .312 .433 .465 .334 .505 .604 .426 .451 .403 .474 .447
[…]
@@
ReProbe, Step-Level (Proposed) 32K .534 .650 .281 .793 .817 .789 .344 .450 .488 .639 .582
ReProbe, Token-Level 32K –.046 –.049 –.072 –.129 –.070 –.017 –.021 –.075 –.055 –.063 –.060
Linear Probe, Token-Level 32K –.112 –.036 –.063 –.157 –.224 –.422 +.011 +.009 –.070 –.157 –.124
Table 9: Replacing the proposed ReProbe with other architectures. Change in PR-AUC relative to the original step-level ReProbe.
@@
Figure 5: Top 2 rows: PR-AUC of ReProbes with increasing training set size (x-axis). Bottom 2 rows: scaling training data either by adding new unique questions or by sampling additional trajectories.
A.4 Impact of Reasoning Length on ReProbe Performance
Figure 6: ROC-AUC on the step-level MATH benchmark using Qwen3-8B, evaluated across bins of reasoning chain length (number of reasoning steps), for the proposed ReProbe and two PRM models.
Figure 6 compares the performance of ReProbe with two selected PRM models across different bins of total reasoning steps. The results indicate that ReProbe consistently outperforms the PRM baselines on shorter generations (i.e., $\leq 8$ reasoning steps), while achieving comparable performance as the reasoning chains become longer.
A.5 Computational Efficiency
To assess computational efficiency, we benchmarked the runtime of ReProbe and several PRM models on 500 test samples from the MATH dataset. The runtime excludes LLM generation and step-extraction overhead. For ReProbe, we measure only the feature-extraction stage together with the forward pass of the ReProbe classifier; for PRMs, we measure solely the inference time of each model. PRM runtimes are based on the official implementations provided in their respective Hugging Face repositories, while ReProbe uses our own implementation of feature extraction and inference. All evaluations were performed with a batch size of 1. Table 10 presents the results.
The results show that ReProbe is by far the fastest method, which is expected given its lightweight architecture and small parameter footprint. In contrast, PRM models incur substantially higher inference costs, often by one to two orders of magnitude. Minor non-monotonicities in runtime across PRMs (e.g., smaller models being slower than larger ones) are attributable to differences in engineering and implementation rather than model size. A theoretical analysis of time and memory complexity can be found in Appendix E.
Method Runtime
[…]
Skywork-PRM-1.5B 17 s
H4-Qwen2.5-PRM-1.5B-0.2 8 min 20 s
[…]
Qwen2.5-Math-7B-PRM800k 37 s
Qwen2.5-Math-PRM-7B 34 s
Math-Shepherd-PRM-7B 5 min 51 s
RLHFlow-PRM-Mistral-Data 5 min 19 s
RLHFlow-PRM-Deepseek-Data 5 min 33 s
Universal-PRM-Qwen2.5-Math-7B 4 min 22 s
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Hidden States, Self-anno (Ours) 14 s
$\bigstar$ ReProbe, Hidden States, DeepSeek-anno (Ours) 13 s
Table 10: Runtime comparison of ReProbe and PRMs on 500 MATH test samples (batch size = 1).
Appendix B ReProbe Training Details
B.1 Training Data Annotation Details
In our experiments, we employ two LLM judges for the annotation of ReProbe training data. For DeepSeek-R1, we follow the officially recommended inference setting, using a temperature of 0.6. Similarly, when using Qwen3-8B as a training data annotator, we also follow the recommended inference hyperparameter, using a temperature of 0.7, a top_k of 20, and a top_p of 0.95. We access Qwen3-8B through vLLM local deployment and access DeepSeek-R1 through the DeepSeek API. The annotation prompts are detailed in Fig 9. For annotation efficiency, we restrict the generation of Qwen3 8B to $256$ tokens during training.
Qwen3-1.7B, with native thinking enabled, generates long CoTs enclosed within <think> tags, without explicit step separations. To train ReProbe for native reasoning, we modify the experimental setup from in two ways: (1) since the native thinking trajectory is much longer than the formatted CoT, we replace the DeepSeek-R1 API with the locally deployed GPT-OSS-120B for step-level correctness annotation to reduce API costs. This model outperforms the January version of DeepSeek-R1 on multiple benchmarks (e.g., GPQA Diamond, AIME 2024). (2) Without format-based step separations, we treat each sentence as a step and prompt GPT-OSS-120B to identify errors conditioned on all previous sentences. For the native thinking mode, during training, we restrict the generation length to $1024$ tokens.
B.2 Training Hyperparameter Details
For all experiments on Phi-4 and Qwen3-8B, we use the same set of hyperparameters. We use a learning rate of 5e-4, a batch size of 128, and a positive class weight of 3. For ReProbes, we use one layer of the transformer encoder with a hidden size of 512 and 16 attention heads. All training continues for 5 epochs. We use 10% of the training data and a 200-sample set of GSM8K, ScienceQA, and StrategyQA for validation and best checkpoint selection.
B.3 Training Cost
All experiments are carried out on cluster nodes with 4 GH200 GPUs and another cluster node with 2 H100 GPUs. Training one ReProbe on 32K datapoints with Hidden States as features takes ~4 GH200 GPU hours, while the variant with Attn+Logit as features takes ~32 GH200 GPU hours. The training of the Attn+Logit ReProbe is slower since extracting attention maps prohibits the use of efficient attention implementations (e.g., Flash Attention).
The total annotation cost for the training dataset (32K samples drawn from PRM800k) was approximately \$200 per model. In addition, around \$300 was spent on evaluation across eight datasets (three for math, three for planning, and two for QA), including the annotation of both step-level correctness and complete reasoning chains for the evaluated decoding strategies.
Appendix C Additional Details of Evaluation Setup
C.1 Baselines
Small PRMs (1.5B): we use 2 PRMs fine-tuned from Qwen2.5-Math-1.5B: Skywork-PRM-1.5B (He et al., 2024) and H4-Qwen2.5-PRM-1.5B-0.2 (HuggingFaceH4, 2025).
Large PRMs (7–8B): we use (1) Math-Shepherd-PRM-7B (Wang et al., 2024), which determines the process labels for each step by estimating the empirical probability of reaching the correct final answer (MC estimation); (2) RLHFlow-PRM-8B-DeepSeek/Mistral models (Xiong et al., 2024) trained with MC-estimated labels from DeepSeek/Mistral rollouts; (3) Universal-PRM-7B (AURORA, 2025) trained using ensemble prompting and reverse verification; (4) Qwen2.5-Math-7B-PRM800k trained on the PRM800k dataset (Lightman et al., 2024); and (5) Qwen2.5-Math-PRM-7B (Zhang et al., 2025c), which combines MC estimation with LLM-as-a-Judge consensus and currently achieves the best result on ProcessBench compared to PRMs of similar scale and computation (Zhao et al., 2025).
UQ methods evaluated in our experiments fall into two categories, differing in computational cost: (1) lightweight scores that use only single generation: Maximum Sequence Probability, Mean Token Entropy, Perplexity (Fadeeva et al., 2023), P(True) (Kadavath et al., 2022), CCP (Fadeeva et al., 2024), and Self-Certainty (Kang et al., 2025), which show significant advantages in reasoning chain selection (Fu et al., 2025); and (2) sampling-based scores: Semantic Entropy, Lexical Similarity (Kuhn et al., 2023), and Degree Matrix (Lin et al., 2024). To compute the latter, we draw $M=10$ alternative steps per step position, making them much less computationally efficient. All methods are implemented using the LM-Polygraph framework (Fadeeva et al., 2023; Vashurin et al., 2025).
C.2 Test Dataset Details
The mathematical domain includes MATH (high school and competition-style math problems; Hendrycks et al., 2021), GSM8K (grade-school math word problems; Cobbe et al., 2021), and ProofNet (natural language proofs of undergraduate-level math problems; Azerbayev et al., 2023). Planning domain includes NaturalPlan (Zheng et al., 2024) (spans three real-world tasks – trip planning, meeting planning, and calendar scheduling). The QA domain includes ScienceQA without multi-modal context (Lu et al., 2022) (covers 26 science subjects from elementary to high school) and StrategyQA (Geva et al., 2021) – a general knowledge QA benchmark that requires implicit multi-step reasoning. Due to computational limitations, we evaluate on randomly sampled subsets of the test sets.
Table 11 presents the statistics of the test datasets used for the step-level benchmark. Importantly, we rely on DeepSeek-R1 for step-level evaluation, which is an expensive API-based LLM. We thus take a subset of test sets for this evaluation. While there are only a few hundred questions, there are a substantial number of reasoning steps, which allows us to draw representative insights.
Dataset Model #Questions #Steps
Mean
#Steps
%Correct
Steps
Answer Mean
Len (Tokens)
MATH Qwen3-8B 200 1203 6.2 85.5% 204
Phi-4 200 2296 11.5 89.2% 417
GSM8k Qwen3-8B 200 1056 5.1 94.0% 139
Phi-4 200 1286 6.4 95.7% 174
ProofNet Qwen3-8B 186 1211 6.5 81.9% 236
Phi-4 186 2227 12.0 91.1% 515
Trips Qwen3-8B 320 5492 17.2 46.9% 641
Phi-4 320 4486 14.0 44.87% 584
Meetings Qwen3-8B 200 2569 12.8 40.9% 739
Phi-4 200 3350 16.75 52.3% 584
Calendar Qwen3-8B 200 1616 8.1 52.5% 365
Phi-4 200 2172 10.86 66.9% 428
StrategyQA Qwen3-8B 500 2865 5.73 81.1% 121
Phi-4 500 3288 6.6 82.6% 182
ScienceQA Qwen3-8B 500 2621 5.24 88.7% 110
Phi-4 500 3415 6.8 91.9% 172
Table 11: Test dataset statistics used for step-level benchmark.
C.3 Test Dataset Examples
Table 12 presents example datapoints randomly picked from our benchmarking datasets.
Dataset Question Answer
- Step 2: Natalia sold half as many clips in May, which is $48\div 2=24$.
- Step 3: Add them: $48+24=72$.
<Answer>: 72
- Step 2: Then x=(r+x)rx=(r+x)-r is rational.
- Step 3: This contradicts $x$ being irrational.
- Step 4: Therefore $r+x$ is irrational.
<Answer>: $r+x$ is irrational
- Step 2: $\sqrt{50}\approx 7.071$.
- Step 3: Sum $\approx 12.548$.
- Step 4: Lies between 12 and 13.
<Answer>: 12 and 13
Trips You plan to visit 3 European cities for 14 days.
Stay in Istanbul for 5 days, Tallinn for 4 days, Zurich for 7 days.
From day 1-7, there is a show in Zurich.
Find a valid trip plan.
- Step 1: Start in Zurich for days 1-7 (show).
- Step 2: Fly Zurich $\to$ Tallinn, stay days 8-11.
- Step 3: Fly Tallinn $\to$ Istanbul, stay days 12-14.
<Answer>: Day 1-7: Zurich, Day 8-11: Tallinn, Day 12-14: Istanbul
Meetings You arrive at Financial District at 9:00AM.
Richard: Marina District 3:30-5:30PM (need 90 min).
Andrew: Alamo Square 5:00-10:00PM (need 120 min).
Travel: FD-Marina 15min, FD-Alamo 17min, Marina-Alamo 15min.
Plan a schedule to meet both.
- Step 1: Leave FD 2:45PM, arrive Marina 3:00PM.
- Step 2: Meet Richard 3:30-5:00PM.
- Step 3: Leave Marina 5:15PM, arrive Alamo 5:30PM.
- Step 4: Meet Andrew 5:30-7:30PM.
<Answer>: FD $\to$ Marina $\to$ Alamo schedule works
Laura busy: 11:30-12:30, 14:30-15:00, 16:00-16:30.
Paul busy: 9:30-10:00, 11:00-14:30, 15:00-17:00.
Paul prefers not after 9:30.
- Step 1: Laura free 9:00-11:30.
- Step 2: Paul free 9:00-9:30 and before 9:30 only.
- Step 3: Intersection is 9:00-9:30.
<Answer>: Meeting 9:00-9:30
- Step 2: Smartphones invented 21st century (iPhone 2007).
- Step 3: Impossible for him to have used one.
<Answer>: No, he did not
ScienceQA Which word is not like the others?
A. horse
B. goat
C. squirrel
D. leg
- Step 1: Horse, goat, squirrel are animals.
- Step 2: Leg is a body part.
- Step 3: Outlier is D.
<Answer>: D. leg
Appendix D Human Validation of LLM-as-a-Judge
DeepSeek-R1’s judgement for step-level correctness is compared against human annotations. The results are shown in Table 13. Besides the 17K PRM800K labels, we manually annotate 1000 steps spanning all studied domains. Here are the sampling details.
We first randomly shuffle the DeepSeek-R1 annotated test set. The human annotation then starts from the first row of the shuffled dataset. Human annotators are not provided with DeepSeek-R1’s rationales (i.e., contents within the <think> tag) and must independently assess the correctness of each reasoning step, following the step-level correctness definition in Fig 9.
The annotation difficulty varies by task. ScienceQA is the easiest, as it mostly requires common sense and usually takes 1 to 2 minutes to annotate one step. StrategyQA requires extensive fact-checking against Wikipedia; however, since the dataset provides relevant facts for grading (available to both human and LLM annotators),annotation usually takes around 2 minutes per step. ProofNet and three planning datasets are the most demanding, often requiring more than 5 minutes per step due to the reasoning length and question complexity.
Dataset # Steps Acc.
PRM800K 17,067 95.29
ProofNet 193 87.05
Trips 102 93.06
Meetings 101 95.70
Calendar 102 91.09
StrQA 313 95.85
SciQA 245 99.28
Appendix E Theoretical Analysis of ReProbe Efficiency
Let the target LLM generate a reasoning trace of total length $T$ tokens, segmented into $S$ reasoning steps with token lengths $\{m_{s}\}_{s=1}^{S}$ such that $\sum_{s=1}^{S}m_{s}=T$, and define $m_{\max}\coloneqq\max_{s}m_{s}$. A PRM is a separate Transformer critic with $L_{\text{prm}}$ layers, hidden width $d_{\text{prm}}$, and parameter count $P_{\text{prm}}$, which scores (possibly growing) text prefixes. With standard dense self-attention, evaluating the PRM over a length-$T$ sequence incurs time at least $\Omega\!\big(L_{\text{prm}}\,d_{\text{prm}}\,T^{2}\big)$ (quadratic attention) and memory $\mathcal{O}\!\big(P_{\text{prm}}+L_{\text{prm}}\,T\,d_{\text{prm}}\big)$ (parameters plus KV-cache).
Hidden-states ReProbe instead taps per-token hidden representations from the frozen target LLM. Let $f$ be the probe’s per-token feature dimension. ReProbe is a lightweight Transformer with $L_{\text{rp}}$ layers, hidden width $d_{\text{rp}}$, and parameter count $P_{\text{rp}}$, applied within each step to the length-$m_{s}$ feature sequence, preceded by a token-wise linear projection $\mathbb{R}^{f}\!\to\!\mathbb{R}^{d_{\text{rp}}}$. ReProbe’s total time complexity ($TC_{rp}$) over the full trace:
$TC_{rp}$ $=\mathcal{O}\!\Big(\sum_{s=1}^{S}\![m_{s}fd_{\text{rp}}+L_{\text{rp}}(m_{s}^{2}d_{\text{rp}}+m_{s}d_{\text{rp}}^{2})]\Big)$ (1)
$=\mathcal{O}\!\Big(Tfd_{\text{rp}}+L_{\text{rp}}(d_{\text{rp}}\!\textstyle\sum_{s}m_{s}^{2}+d_{\text{rp}}^{2}T)\Big).$
where $m_{s}fd_{\text{rp}}$ is the cost of the per-token feature projection, $m_{s}^{2}d_{\text{rp}}$ is the quadratic self-attention mixing cost within a step, and $m_{s}d_{\text{rp}}^{2}$ accounts for the token-wise dense linear maps inside each Transformer block (e.g., Q/K/V and output projections, as well as the MLP). Its memory is $\mathcal{O}\!\big(P_{\text{rp}}+m_{\max}(f+L_{\text{rp}}d_{\text{rp}})\big)$ when step-local features/KV-cache are discarded after scoring each step. Using $\sum_{s}m_{s}^{2}\leq m_{\max}T$, we obtain the upper bound
$TC_{rp}\leq\mathcal{O}\!\left(Tfd_{\text{rp}}+L_{\text{rp}}(d_{\text{rp}}m_{\max}T+d_{\text{rp}}^{2}T)\right),$
which is near-linear in $T$ whenever $m_{\max}\ll T$ (equivalently, $S$ grows with $T$). By contrast, PRM-based verification remains at least quadratic in $T$ due to attention over length-$T$ prefixes and additionally incurs substantially larger memory complexity (typically $P_{\text{rp}}\!\ll\!P_{\text{prm}}$).
Appendix F Prompts
F.1 Prompts for Reasoning